te 24 teet bung nd え나
Answer #1

Lets take a horizontal strip of width dh at depth h .

then length of that strip will be , L = 24 - [ (24-15)h/12 ]

L = 24 - 2h + 1.25h = 24 - 0.75h

Pressure at depth h, P = rho*g*h

Area = L dh = (24 - 0.75h)dh


Force of that strip = Pressure x area = rho*g*h(24 - 0.75h)dh

dF = rho*g*h(24 - 0.75h)dh .........Ans

where rho is density of water.

g = acc due to gravity

and h is from 0 to 12 feet.


F = rho*g [ 12h^2 - 0.25 h^3 ]